OpenStudy (anonymous):

Solve for x: 3|x - 3| + 2 = 14

OpenStudy (anonymous):

what do you think you should do first?

OpenStudy (anonymous):

the lx-3l part?

OpenStudy (jdoe0001):

\(\bf 3|x - 3| + 2 = 14\implies 3|x - 3|=12\implies |x - 3|=4\\ \quad \\ |x - 3|=4\implies \begin{cases} +(x - 3)=4\\ \quad \\ \bf -(x - 3)=4 \end{cases}\)

OpenStudy (anonymous):

I still don't understand what the answer would be

OpenStudy (anonymous):

No solutions?

OpenStudy (♪chibiterasu):

It would be 1 and 7.

OpenStudy (anonymous):

Okay..

OpenStudy (anonymous):

-1 and 7?

OpenStudy (♪chibiterasu):

No, 1 and 7.

OpenStudy (anonymous):

Thats not a possible answer...

OpenStudy (♪chibiterasu):

lol @jdoe0001

OpenStudy (anonymous):

The answers can be No solutions x = -1, x = 8.3 x = 0, x = 7 x = -1, x = 7

OpenStudy (♪chibiterasu):

I guess C then :I

OpenStudy (anonymous):

lol

OpenStudy (jdoe0001):

hmm SaviClouse so what did you get for each scenario?

OpenStudy (anonymous):

Huh?

OpenStudy (jdoe0001):

do you even know ... you're supposed to solve 2 equations?

OpenStudy (anonymous):

Yeah and the answer is C right?

OpenStudy (jdoe0001):

dunno... is that what you got?

OpenStudy (anonymous):

Yeah but I don't think it's right